OpenStudy (anonymous):

Jessica deposits $300 into a savings account that pays a simple interest rate of 60%. How much money will Jessica have in her account at the end of six months ?

OpenStudy (anonymous):

Simple Interest =P*R*T/100 P=principal = 300 R=rate of interest=60% T=Time=6 months= 1/2year 300*60*1/2*100=9 In the account she will be having P+9$=300+9=309$

OpenStudy (anonymous):

wow 60% interest rate